OpenStudy (anonymous):

a=3n+1 solve for n i need a step by step simple explanation of how to do this problem

OpenStudy (anonymous):

subtract the one to the other side and your left with a-1=3n divide by 3 (a-1)/3=n thats your answer ^^

OpenStudy (anonymous):

We have two unknown variable so all we can do is isolate n. \[\large{a=3n+1}\]Subtract both sides by one.\[\large{a - 1=3n}\]Divide both sides by three.\[\boxed{\large{\dfrac{a - 1}{3}=n}}\]

OpenStudy (anonymous):

can you explain how you got the answer

OpenStudy (anonymous):

We already explained our answer. What don't you understand?